OpenStudy (anonymous):

(square root of 75b over c) plus (square root of 4b over 3c) plus (square root of 25b over c),

OpenStudy (anonymous):

pls show the complete solution tnx.

OpenStudy (anonymous):

\[\sqrt{\frac{75b}{c}}+\sqrt{\frac{4b}{3c}}+\sqrt{\frac{25b}{c}}\] like this?

OpenStudy (anonymous):

or is the c outside of the radical?

OpenStudy (anonymous):

yes the first one